OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9>

OpenStudy (anonymous):

the answer choices are; 81.6° 25.8° 35.8° 71.6°

OpenStudy (amistre64):

youll need to know the dot product, and the length of each vector ... \[cos(\alpha)=\frac{u\cdot v}{|u||v|}\]